# ECG-Byte: A Tokenizer for End-to-End Generative Electrocardiogram Language Modeling

### v1

@@ 5.3 Efficiency of ECG-Byte
tab:cross Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1 $L_{CL}$ ECG-QA MIMIC-IV ECG-QA PTB-XL 11.64 ± 0.45 41.48 ± 0.11 25.74 ± 0.13 91.24 ± 0.05 $L_{MIM}$ 11.70 ± 0.29 42.22 ± 0.28 26.41 ± 0.10 91.51 ± 0.03 $L_{MERL}$ (Liu et al., 2024) 11.53 ± 0.19 39.23 ± 0.40 25.58 ± 0.28 91.59 ± 0.03 $L_{CL}$ + $L_{MIM}$ 9.71 ± 0.10 35.10 ± 0.28 24.91 ± 0.19 87.88 ± 0.08 ECG-Byte 8.70 ± 0.04 40.39 ± 0.40 23.29 ± 0.18 91.51 ± 0.03 $L_{CL}$ ECG-QA PTB-XL ECG-QA MIMIC-IV 5.10 ± 0.04 22.77 ± 0.28 14.63 ± 0.32 77.89 ± 0.13 $L_{MIM}$ 7.68 ± 0.46 35.77 ± 0.13 22.32 ± 0.33 90.28 ± 0.07 $L_{MERL}$ (Liu et al., 2024) 7.39 ± 0.15 28.33 ± 0.58 18.59 ± 0.35 89.30 ± 0.05 $L_{CL}$ + $L_{MIM}$ 7.49 ± 0.21 30.53 ± 0.59 20.25 ± 0.27 86.53 ± 0.11 ECG-Byte 7.86 ± 0.13 35.01 ± 0.41 21.49 ± 0.24 90.29 ± 0.07
[…]
tab:efficiency 1st Stage 2nd Stage ECG-Byte end-to-end # of Data 2,513,435 400,000 1,000,000 400,000 Total # of Data 2,913,435 1,400,000 Time (minutes) $\sim$1258.50 $\sim$469.25 $\sim$385.12 $\sim$420.32 Total Time (minutes) $\sim$1727.75 $\sim$805.44
Table 3: Efficiency of our method compared against 2-stage pretraining methods.
5.4 Ablation Study
We conduct several ablation studies to show the variability of performance with ECG-Byte when we alter the LLM used for finetuning, use different sequence lengths $L$ when inputting to the LLM, training ECG-Byte with various number of merges num_merges, and varying ECG lengths $T$. With the exception of the ablating parameter, we fix all other parameters to num_merges = 3500, $L=1024$, and $T=500$. We report results on the test set of the PTB-XL variant of ECG-QA (Oh et al., 2023) unless specified otherwise.
[…]
tab:diff_llm LLM BLEU-4 Rouge-L Meteor BertScore F1 GPT2 XL 1.5B (Radford et al., 2019) 12.30 ± 0.19 41.33 ± 0.57 26.48 ± 0.33 92.00 ± 0.06 Gemma 2B (Team et al., 2024) 13.78 ± 0.18 45.48 ± 0.55 28.32 ± 0.23 92.01 ± 0.02 OPT 1.3B (Zhang et al., 2022) 12.26 ± 0.20 41.84 ± 0.52 26.21 ± 0.29 91.78 ± 0.04 Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Table 4: Ablation study on using different LLMs.
Different LLMs
We show the variability in performance of ECG-Byte when using different LLMs with similar numbers of parameters in Table LABEL:tab:diff_llm. While Llama 3.2 1B (Grattafiori et al., 2024) achieves the best results, GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022) also deliver comparable performances. These findings demonstrate that our method is not limited to Llama 3.2 1B but can achieve similar results across a variety of LLMs.
Sequence Length
Input lengths for LLMs are an important parameter to consider for efficient training since the calculation of attention is quadratic with respect to the input length (Vaswani et al., 2023). We present results on different sequence lengths $L$ in Table LABEL:tab:seq_len. Although the difference is not substantial, we can see that when $L=1024$ and $L=2048$ the model yields higher performance then $L=512$. We attribute this to the rate of truncation and padding for the encoded ECG. Observing Figure 2, most ECGs were being encoded to token sequence lengths of around 500 to 1500. Therefore, we hypothesize that when $L=512$ a large portion of the ECG token sequence gets truncated, resulting in lower performance.
[…]
tab:seq_len $L$ BLEU-4 Rouge-L Meteor BertScore F1 512 13.61 ± 0.15 48.15 ± 0.57 29.10 ± 0.28 92.41 ± 0.05 1024 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 2048 13.88 ± 0.22 45.21 ± 0.48 28.31 ± 0.27 90.88 ± 0.02
Table 5: Ablation study on varying sequence lengths $L$.
Number of Merges
The number of merges num_merges performed during training ECG-Byte corresponds to how much the algorithm compresses the concatenated sequence of quantized ECGS $\mathbf{X}_{\text{concat}}$. More num_merges means more compression, which can affect the expressiveness of the encoded sequence. In Table LABEL:tab:num_merges, we show the performance of our method with different num_merges. The results indicate that while performance varies slightly with the number of merges, values of num_merges greater than 500 generally yield similar outcomes.
[…]
tab:num_merges num_merges BLEU-4 Rouge-L Meteor BertScore F1 500 13.61 ± 0.53 46.50 ± 0.28 28.49 ± 0.49 92.33 ± 0.02 1750 14.50 ± 0.25 46.74 ± 0.48 30.03 ± 0.25 92.55 ± 0.01 2500 15.10 ± 0.39 46.37 ± 0.28 30.12 ± 0.23 92.53 ± 0.05 3500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Table 6: Ablation study on varying number of merges num_merges.
@@ 5.4 Ablation Study
Lastly, we show the effect of the length $T$ being considered when encoding the ECG with ECG-Byte in Table LABEL:tab:len_ecg. We want to note that for the results of $T=2000$, the full unsegmented ECG is utilized. Consequently, the number of instances available is less than the targeted dataset size of 400,000 (i.e., 97,244). Thus, when $T=2000$, we use the full dataset to train the model. For shorter segment lengths, such as $T=250$ and $T=500$, the model demonstrates strong performances indicating that shorter segments can effectively preserve relevant information for NLG. Interestingly, for $T=2000$, the model achieves the highest performance across all metrics. This suggests that when the model is trained with the full 10 second encoded ECG, it benefits from richer contextual information present in the complete ECG waveform.
[…]
tab:len_ecg $T$ BLEU-4 Rouge-L Meteor BertScore F1 250 12.64 ± 0.20 47.31 ± 0.26 27.97 ± 0.21 92.32 ± 0.06 500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 1250 11.01 ± 0.19 43.84 ± 0.28 25.49 ± 0.20 93.07 ± 0.03 2000 14.54 ± 0.17 48.03 ± 0.27 32.11 ± 0.22 92.91 ± 0.04
Table 7: Ablation study on varying lengths $T$.
@@ 5.5 ECG-Byte Analysis
We examine the token usage and length distributions for ECG-Byte with num_merges = 3500 on a subsample of 277,840 ECGs from the PTB-XL dataset. The left panel of Figure 2 displays the token usage distribution, showing token frequency (y-axis) ranked in descending order (x-axis). A small subset of tokens dominates the occurrences, while the rest are infrequently used—a typical characteristic of BPE-based tokenization, where common patterns are compressed into frequent tokens and rare patterns into infrequent ones. The right panel of Figure 2 illustrates the token length distribution of the encoded ECGs, with most falling between 500 and 1000 tokens, demonstrating ECG-Byte’s effective compression of the original signal. Additional examples of these distributions are provided in Appendix A.3.
Token to ECG Mapping
Figure 3: A mapping between tokens used for a given ECG Lead I. More examples are provided in Appendix A.2.
To illustrate how ECG-Byte encodes ECG signals, we analyze the mapping between tokens and signal features. Figure 3 shows an example Lead I ECG signal with unique token IDs (represented by different colors) overlayed. The P wave, QRS complex, and T wave are distinctly captured by different tokens, though this precision varies across instances. As demonstrated, ECG-Byte effectively merges key regions of the signal. Additional examples are provided in Appendix A.2 due to page limitations.
Attention Visualizations
Figure 4 visualizes attention weights across a selected ECG lead and text portions of the input after training. We focus on one lead due to the uniformity of attention across encoded signal tokens. For interpretability, the reversed ECG signal is overlayed on the encoded ECG. The model primarily attends to the textual portion of the input sequence, as shown in Figure 4. Previous studies have debated whether attention visualizations are inherently explainable (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019) and explored their role in vision-language models (Aflalo et al., 2022; Woo et al., 2024; Arif et al., 2024; Cui et al., 2024). These works often observe minimal attention to visual input, with models relying primarily on text. We hypothesize that a similar phenomenon occurs in Figure 4, as the ECG tokens, though represented like text, are 1) newly introduced and 2) perceived as a different modality (e.g., vision). Additional examples are provided in Appendix A.4.
6 Discussion and Conclusion
In this study, we introduce ECG-Byte, a custom BPE algorithm to encode ECGs into a discrete sequence of tokens for conditional autoregressive NLG. ECG-Byte introduces a paradigm shift in generative ECG language modeling by enabling efficient end-to-end training, compared to traditional two-stage pretraining approaches. Our pipeline demonstrates strong performances, achieving results comparable to two-stage methods while requiring only half the training time and approximately $48\%$ of the data. In addition to its efficiency, ECG-Byte enhances interpretability.
Figure 4: The attention weight overlayed on top of both the text (top) and ECG (bottom). More examples are provided in Appendix A.4.
By analyzing its underlying mechanism, we observe that critical ECG regions, such as the P wave, the QRS complex, and the T wave, are effectively grouped during tokenization, as illustrated in Figure 3. Furthermore, the reversibility of the compressed token sequence allows us to trace each token back to its original ECG signal segment, providing insight into the specific portions of the signal attended to by the model. However, as shown in Figure 4, the model’s attention weight distribution resembles that of vision language models, focusing primarily on the textual components of the input sequence during generation.
This work is in its early stages and needs further exploration. Future directions include: (1) refining BPE merging rules to better capture ECG-specific features, (2) adopting more advanced quantization techniques that preserve time-series characteristics (Carson et al., 2024b; Elsworth and Güttel, 2020; Carson et al., 2024a), (3) introducing stronger modality-specific distinctions, such as embeddings beyond [SIG_START] and [SIG_END] (Gui et al., 2023), and (4) extending ECG-Byte for conversational tasks through instruction tuning.
Limitations
One of the main limitations of this work is the scale in terms of computing and data. Since we only used a subset of the data to train and test our method, we were unable to train the model to its full potential. However, even with only using a small subset of the data, we are able to see extremely promising results compared to other 2-stage SSL pretraining methods. Therefore, we do not view this limitation as a major bottleneck.
[…]
@@ 6 Discussion and Conclusion
This work is done in collaboration with the Mario Lemieux Center for Heart Rhythm Care at Allegheny General Hospital. We thank Wenhao Ding, Haohong Lin, Shiqi Liu, and Hyoeun Kang for the valuable discussions.
Appendix A ECG-Byte
A.1 Additional Pseudocode for ECG-Byte
Input: Array of IDs ids, Pair to merge pair as $(u_{1},u_{2})$, New ID new_id
Output: Merged vector of IDs
Function merge(ids, pair, new_id):
       Initialize empty vector new_ids with capacity of ids;
       i $\leftarrow$ 0;
       while i $<$ length of ids do
[…]
                   new_ids.append(new_id);
                   i $\leftarrow\texttt{i}+2$;
             end if
            else
                   new_ids.append(ids[i]);
                   i $\leftarrow\texttt{i}+1$;
             end if
       end while
      return new_ids;
Algorithm 3 Merging a pair in an ID array
Input: Array of IDs ids
Output: HashMap of pairs and their frequencies
Function get_stats(ids):
       pair_counts $\leftarrow$ Parallel fold operation:;
       foreach window of size 2 in ids do
             Let $(u_{1},u_{2})\leftarrow$ elements of the window;
             Increment the count of $(u_{1},u_{2})$ in local pair_count;
       end foreach
      pair_counts $\leftarrow$ Parallel reduce operation to combine local pair_count HashMaps;
       return pair_counts;
Algorithm 4 Calculating Frequency of Byte Pairs in an Array
A.2 Mapping between Token and ECG
@@
We add more examples of the mapping between the ECG signal and the encoded tokens for ECG-Byte in Figures 5 and 6.
Figure 5: A mapping between tokens used for a given ECG Leads I, II, III, aVL, aVR, aVF.
Figure 6: A mapping between tokens used for a given ECG Leads V1, V2, V3, V4, V5, V6.
A.3 Token usage and length distribution for varying num_merges
@@
A.4 Attention Visualizations
We add more visualizations of the attention weights in Figures 8, 9, 10, 11, 12, 13, 14, 15.
Figure 8: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 9: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 10: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 11: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 12: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 13: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 14: The attention weight overlayed on both text (top) and ECG (bottom).
Figure 15: The attention weight overlayed on both text (top) and ECG (bottom).
Appendix B 2-stage Pretraining Approaches
@@ A.4 Attention Visualizations
B.1 Contrastive learning approaches
@@
$\mathcal{L}_{\text{CL}}=-\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}$ $\log\frac{\exp\left(\text{sim}(z_{i}^{\text{img}},z_{i}^{\text{txt}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{img}},z_{j}^{\text{txt}})/\tau\right)}$
$+$ $\log\frac{\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{i}^{\text{img}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{j}^{\text{img}})/\tau\right)}\Bigg{]}$
@@
B.2 Masked image modeling approaches
@@
B.3 Dual approaches
@@
Appendix C Additional Results
C.1 Does Larger LLMs Yield Higher Performance?
We present the results of ablating the size of the LLM in Table LABEL:tab:large_llm. Interestingly, the performance across the three different model sizes (1B, 3B, 8B) remains fairly similar. We believe that the limited dataset size prevents the larger models from realizing their full performance potential. We hypothesize that increasing the amount of training data would enable the larger models to leverage their greater capacity, resulting in observable performance improvements.
@@ C.1 Does Larger LLMs Yield Higher Performance?
tab:large_llm LLM BLEU-4 Rouge-L Meteor BertScore F1 Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07 Llama 3.2 3B (Grattafiori et al., 2024) 14.80 ± 0.17 46.55 ± 0.21 29.53 ± 0.16 92.42 ± 0.01 Llama 3.1 8B (Grattafiori et al., 2024) 13.80 ± 0.16 46.29 ± 0.25 28.56 ± 0.11 92.44 ± 0.05

### v2

@@ 5.4 Ablation Study
Lastly, we show the effect of the length $T$ being considered when encoding the ECG with ECG-Byte in Table 6. We want to note that for the results of $T=2500$, the full unsegmented ECG is utilized. Consequently, the number of instances available is less than the targeted dataset size of 400,000 (i.e., 97,244). Thus, when $T=2500$, we use the full dataset to train the model. For shorter segment lengths, such as $T=250$ and $T=500$, the model demonstrates strong performances indicating that shorter segments can effectively preserve relevant information for NLG. Interestingly, for $T=2500$, the model achieves the highest performance across all metrics. This suggests that when the model is trained with the full 10 second encoded ECG, it benefits from richer contextual information present in the complete ECG waveform.
@@ 5.5 ECG-Byte Analysis
We examine the token usage and length distributions for ECG-Byte with num_merges = 3500 on a subsample of 277,840 ECGs from the PTB-XL dataset. For a 2-second ECG ($T=500$), the raw signal comprises $12\times 500=6000$ symbols. ECG-Byte then compresses these 6000 symbols at an average ratio of 12.66x, yielding approximately 500 tokens per ECG on average, as illustrated in Figure 2. The left panel of Figure 2 displays the token usage distribution, showing token frequency (y-axis) ranked in descending order (x-axis). A small subset of tokens dominates the occurrences, while the rest are infrequently used—a typical characteristic of BPE-based tokenization, where common patterns are compressed into frequent tokens and rare patterns into infrequent ones. The right panel of Figure 2 illustrates the token length distribution of the encoded ECGs, with most falling between 500 and 1000 tokens, demonstrating ECG-Byte’s effective compression of the original signal. Additional examples of these distributions are provided in Appendix B.3.
Figure 3: A mapping between tokens used for a given ECG Lead I. More examples are provided in Appendix B.2.
Token to ECG Mapping
To illustrate how ECG-Byte encodes ECG signals, we analyze the mapping between tokens and signal features. Figure 3 shows an example Lead I ECG signal with unique token IDs (represented by different colors) overlaid. The P wave, QRS complex, and T wave are distinctly captured by different tokens, though this precision varies across instances. As demonstrated, ECG-Byte effectively merges key regions of the signal. Additional examples are provided in Appendix B.2 due to page limitations.
Figure 4: The attention weight overlaid on top of both the text (top) and ECG (bottom). More examples are provided in Appendix B.4.
Attention Visualizations
Figure 4 visualizes attention weights across a selected ECG lead and text portions of the input after training. We focus on one lead due to the uniformity of attention across encoded signal tokens. For interpretability, the reversed ECG signal is overlaid on the encoded ECG. The model primarily attends to the textual portion of the input sequence, as shown in Figure 4. Previous studies have debated whether attention visualizations are inherently explainable (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019) and explored their role in vision-language models (Aflalo et al., 2022; Woo et al., 2024; Arif et al., 2024; Cui et al., 2024). These works often observe minimal attention to visual input, with models relying primarily on text. We hypothesize that a similar phenomenon occurs in Figure 4, as the ECG tokens, though represented like text, are 1) newly introduced and 2) perceived as a different modality (e.g., vision). We note that attention visualizations may not inherently indicate which parts of the input sequence contribute to the final generated output. Although methods such as LIME (Ribeiro et al., 2016) or Integrated Gradients (Sundararajan et al., 2017) may provide better attributions, we leave this investigation for future work. Additional examples are provided in Appendix B.4.
6 Discussion and Conclusion
In this study, we introduce ECG-Byte, a custom BPE algorithm to encode ECGs into a discrete sequence of tokens for conditional autoregressive NLG. ECG-Byte introduces a paradigm shift in generative ECG language modeling by enabling efficient end-to-end training, compared to traditional two-stage pretraining approaches. Our pipeline demonstrates strong performance, achieving results comparable to two-stage methods while being about 3 times faster in total training time and requiring approximately 48% of the data. Drawing inspiration from MEIT (Wan et al., 2024), we simulate two-stage pretraining for NLG in an end-to-end manner. Our findings indicate that the initial training stage is essential for achieving performance gains. In addition to its efficiency, ECG-Byte enhances interpretability. By analyzing its underlying mechanism, we observe that critical ECG regions, such as the P wave, the QRS complex, and the T wave, are effectively grouped during tokenization, as illustrated in Figure 3. Furthermore, the reversibility of the compressed token sequence allows us to trace each token back to its original ECG signal segment, providing insight into the specific portions of the signal attended to by the model. However, as shown in Figure 4, the model’s attention weight distribution resembles that of vision language models, focusing primarily on the textual components of the input sequence during generation.
We emphasize that the goal of ECG-Byte is not to claim it as the best method for representing ECG data or training generative, autoregressive Electrocardiogram-Language Models (ELMs). Instead, ECG-Byte demonstrates the potential of using a rule-based compressor rather than a learnable one for ECG data in the context of ELMs. As stated earlier, we find ECG-Byte both highly efficient and intuitive in compressing ECG data, and its tokenization is interpretable because it can be reversed to recover the original signal. We hope the broader research community will recognize the potential of rule-based compressors for generative tasks with different health signal data beyond ECGs. This early-stage work invites the community to contribute to advancing generative ECG language modeling by highlighting some future directions. Future directions include: (1) a comprehensive benchmark with different input representations and training methods for ELMs, (2) refining BPE merging rules to better capture ECG-specific features, (3) adopting more advanced quantization techniques that preserve time-series characteristics (Carson et al., 2024b; Elsworth and Güttel, 2020; Carson et al., 2024a), and (4) introducing stronger modality-specific distinctions, such as embeddings beyond [SIG_START] and [SIG_END] (Gui et al., 2023).
Limitations
One key limitation of this work is the scale of both computing resources and data. Due to limited computing capacity, we trained on only a data subset using a batch size of 2 for a single epoch, which may have constrained the model’s full potential. Nevertheless, as prior work indicates that finetuned LLMs can perform satisfactorily with little data (Brown et al., 2020), the promising results of ECG-Byte—in comparison with other two-stage SSL pretraining methods—suggest that this limitation is not a critical bottleneck.
@@ 6 Discussion and Conclusion
This work is done in collaboration with the Mario Lemieux Center for Heart Rhythm Care at Allegheny General Hospital. We thank Wenhao Ding, Haohong Lin, Shiqi Liu, Hyoeun Kang, and Tony Chen for the valuable discussions.
Appendix A Training Details
A.1 Learning Rate Scheduler
We utilize a custom learning rate scheduler in all of our experiments. This scheduler applies an initial learning rate $\mathrm{init\_lr}$ scaled by the model’s hidden dimension ($\mathrm{d_{model}}^{-0.5}$) and dynamically adjusts it based on training steps, with a warm-up phase of 500 steps. The learning rate at step $\mathrm{n_{steps}}$ is updated as $\text{lr}=\mathrm{init\_lr}\times\min\left(\mathrm{n_{steps}}^{-0.5},\ \mathrm{n_{warmup}}^{-1.5}\times\mathrm{n_{steps}}\right)$.
Appendix B Additional Details Regarding ECG-Byte
B.1 Additional Pseudocode for ECG-Byte
We provide detailed pseudocode for the ECG-Byte encoding process, merge and get_stats functions in Algorithms 4, 2, and 3 respectively.
B.2 Mapping between Token and ECG
@@
We add more examples of the mapping between the ECG signal and the encoded tokens for ECG-Byte in Figure 6.
B.3 Token usage and length distribution for varying num_merges
@@
B.4 Attention Visualizations
We add more visualizations of the attention weights in Figure 7.
Appendix C Two-stage Pretraining Approaches
@@
C.1 Contrastive learning approaches
@@
$\mathcal{L}_{\text{CL}}=-\frac{1}{N}\sum_{i=1}^{N}\Biggl{[}\log\frac{\exp\left(\text{sim}(z_{i}^{\text{img}},z_{i}^{\text{txt}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{img}},z_{j}^{\text{txt}})/\tau\right)}+\log\frac{\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{i}^{\text{img}})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(z_{i}^{\text{txt}},z_{j}^{\text{img}})/\tau\right)}\Biggr{]}$
@@
C.2 Masked image modeling approaches
@@
Algorithm 2 Merging a Pair in an ID Array merge
Input: Array of IDs ids, pair to merge pair as $(u_{1},u_{2})$, new ID new_id.
Output: Updated array of IDs ids with merged pairs.
1:  $i\leftarrow 0$, $write\leftarrow 0$
2:  while $i<\text{len({ids})}$ do
3:     if $i+1<\text{len({ids})}$ and $(\texttt{ids}[i],\texttt{ids}[i+1])=\texttt{pair}$ then
[…]
5:        $write\leftarrow write+1$
6:        $i\leftarrow i+2$
7:     else
8:        Set $\texttt{ids}[write]\leftarrow\texttt{ids}[i]$
9:        $write\leftarrow write+1$
10:        $i\leftarrow i+1$
11:     end if
12:  end while
13:  Truncate ids to length $write$
14:  return ids
C.3 Dual approaches
@@
Algorithm 3 Calculating Frequency of Byte Pairs in an Array get_stats
Input: Array of IDs ids.
Output: HashMap of byte pairs and their frequencies pair_counts.
@@
2:     pair_counts $\leftarrow$ Empty HashMap
3:     for all each window of size 2 in ids do
4:        $(u_{1},u_{2})\leftarrow$ two elements of the window
5:        Increment count for $(u_{1},u_{2})$ in pair_counts
6:     end for
7:  else
8:     pair_counts $\leftarrow$ Parallel fold operation:
9:     for all each window of size 2 in ids (in parallel) do
10:        $(u_{1},u_{2})\leftarrow$ two elements of the window
11:        Increment count for $(u_{1},u_{2})$ in the local HashMap
12:     end for
13:     Combine local HashMaps using a parallel reduce operation to obtain pair_counts
14:  end if
15:  return pair_counts
Appendix D Additional Results and Discussions
Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1
$L_{CL}$ ECG-QA MIMIC-IV ECG-QA PTB-XL 11.64 ± 0.45 41.48 ± 0.11 25.74 ± 0.13 91.24 ± 0.05
$L_{MIM}$ 11.70 ± 0.29 42.22 ± 0.28 26.41 ± 0.10 91.51 ± 0.03
$L_{MERL}$ (Liu et al., 2024a) 11.53 ± 0.19 39.23 ± 0.40 25.58 ± 0.28 91.59 ± 0.03
$L_{Dual}$ 9.71 ± 0.10 35.10 ± 0.28 24.91 ± 0.19 87.88 ± 0.08
ECG-Byte 8.70 ± 0.04 40.39 ± 0.40 23.29 ± 0.18 91.51 ± 0.03
$L_{CL}$ ECG-QA PTB-XL ECG-QA MIMIC-IV 5.10 ± 0.04 22.77 ± 0.28 14.63 ± 0.32 77.89 ± 0.13
$L_{MIM}$ 7.68 ± 0.46 35.77 ± 0.13 22.32 ± 0.33 90.28 ± 0.07
$L_{MERL}$ (Liu et al., 2024a) 7.39 ± 0.15 28.33 ± 0.58 18.59 ± 0.35 89.30 ± 0.05
$L_{Dual}$ 7.49 ± 0.21 30.53 ± 0.59 20.25 ± 0.27 86.53 ± 0.11
ECG-Byte 7.86 ± 0.13 35.01 ± 0.41 21.49 ± 0.24 90.29 ± 0.07
D.1 Cross Dataset Transferability
We present the results of cross-dataset transferability in Table 7, comparing our approach, ECG-Byte, with two-stage pretraining methods. ECG-Byte achieves the best zero-shot transfer performance in BLEU-4 and BertScore F1 scores when transferring from the ECG-QA PTB-XL dataset to the ECG-QA MIMIC-IV dataset. When transferring from the ECG-QA MIMIC-IV dataset to the ECG-QA PTB-XL dataset, although other two-stage pretraining methods demonstrate higher performance, ECG-Byte maintains competitive results across all metrics.
Table 8: Ablation study on varying number of merges num_merges.
num_merges BLEU-4 Rouge-L Meteor BertScore F1
500 13.61 ± 0.53 46.50 ± 0.28 28.49 ± 0.49 92.33 ± 0.02
1750 14.50 ± 0.25 46.74 ± 0.48 30.03 ± 0.25 92.55 ± 0.01
2500 15.10 ± 0.39 46.37 ± 0.28 30.12 ± 0.23 92.53 ± 0.05
3500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Number of Merges
The number of merges num_merges performed during training ECG-Byte corresponds to how much the algorithm compresses the concatenated sequence of quantized ECGS $\mathbf{X}_{\text{concat}}$. More num_merges means more compression, which can affect the expressiveness of the encoded sequence. In Table 8, we show the performance of our method with different num_merges. The results suggest that, although performance fluctuates slightly with the number of merges, varying num_merges generally produces comparable outcomes.
D.2 Does Larger LLMs Yield Higher Performance?
We present the results of ablating the size of the LLM in Table 9. Interestingly, the performance across the three different model sizes (1B, 3B, 8B) remains fairly similar. We believe that the limited dataset size prevents the larger models from realizing their full performance potential. We hypothesize that increasing the amount of training data would enable the larger models to leverage their greater capacity.